OpenStudy (anonymous):

You have 1200 feet of fencing and you want to make two fenced in enclosures by splitting one enclosure in half. What are the largest dimensions of this enclosure that you could build? Answer 125 ft x 200 ft 100 ft x 105 ft 300 ft x 200 ft 105 ft x 120 ft

OpenStudy (anonymous):

@Umangiasd

OpenStudy (anonymous):

@Umangiasd Help please?

OpenStudy (anonymous):

I got A @Umangiasd

OpenStudy (anonymous):

I'm reading, calm xD

OpenStudy (anonymous):

Hahaha Alright.

OpenStudy (anonymous):

I have to assume that this enclosures are quadrilateral?

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

I changed my answer to C but not sure

OpenStudy (anonymous):

yup 300ft and 200 ft. Do you know basic calculus?

OpenStudy (anonymous):

Kinda. Why?

OpenStudy (anonymous):

You can maximize (or minimize) functions by deriving;  Now, your 1200 ft of fence will be reparted between a and b like this: (1) 2b+3a=1200 -->b=(1200-3a)/2 Then, the Area of a quadrilateral is given by the product of its sides, the we have (2) a*b=A, where A=area Now we replace b from (1) in (2) a(1200-3a)/2=A We derive and we have (1/2)*(1200-6a)=dA, we equal dA to 0 in order to maximize and.. a= 200, then because of (1), b= 300